# Tight Robustness Certificates and Wasserstein Distributional Attacks for Deep Neural Networks

### v1

@@ Abstract
Wasserstein distributionally robust optimization (WDRO) provides a framework for adversarial robustness, yet existing methods based on global Lipschitz continuity or strong duality often yield loose upper bounds or require prohibitive computation. In this work, we address these limitations by introducing a primal approach and adopting a notion of exact Lipschitz certificate to tighten this upper bound of WDRO. In addition, we propose a novel Wasserstein distributional attack (WDA) that directly constructs a candidate for the worst-case distribution. Compared to existing point-wise attack and its variants, our WDA offers greater flexibility in the number and location of attack points. In particular, by leveraging the piecewise-affine structure of ReLU networks on their activation cells, our approach results in an exact tractable characterization of the corresponding WDRO problem. Extensive evaluations demonstrate that our method achieves competitive robust accuracy against state-of-the-art baselines while offering tighter certificates than existing methods. Our code is available at https://github.com/OLab-Repo/WDA
@@ 1 Introduction
Modern deep networks achieve remarkable accuracy yet remain fragile to distribution shift and adversarial perturbations (Szegedy et al., 2014; Goodfellow et al., 2014; Kurakin et al., 2018; Hendrycks & Dietterich, 2019; Ovadia et al., 2019; Taori et al., 2020; Koh et al., 2021), raising concerns about their reliability in deployment. A principled avenue for robustness is Wasserstein distributionally robust optimization (WDRO, Mohajerin Esfahani & Kuhn 2018; Gao & Kleywegt 2023), which controls worst-case test risk over an ambiguity set within a Wasserstein ball around the empirical distribution and admits tight dual characterizations from optimal transport (Villani, 2008; Santambrogio, 2015). While numerous defenses have been proposed, a fundamental gap persists between theoretical robustness certificates and practical adversarial evaluation: existing Lipschitz-based certificates often provide loose upper bounds that vastly overestimate the true worst-case loss (Virmaux & Scaman, 2018), while standard attacks restrict perturbations to fixed-radius balls around individual points (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018). This mismatch stems from two limitations: certificates typically rely on global worst-case analysis that ignores the actual network geometry traversed by data, and attacks consider only point-wise perturbations rather than distributional shifts permitted by Wasserstein threat models (Singh et al., 2018; Gao & Kleywegt, 2023). The discrepancy is particularly pronounced for modern architectures with ReLU activations, where the network behaves as a piecewise-affine function whose local properties vary dramatically across regions (Jordan & Dimakis, 2020), and those with smooth activations (GELU, SiLU/Swish) exhibit complex nonlinear geometry (Hendrycks & Gimpel, 2016; Ramachandran et al., 2017; Elfwing et al., 2018). In this work, we aim to address both sides of this gap: our contributions can be summarized as follows.
@@
    For a class of networks with Rectified Linear Unit (ReLU) activations (Nair & Hinton, 2010), we analyze the upper and lower bounds of the Wasserstein Distributional Robust Optimization (WDRO) problem by connecting with the tight Lipschitz constant studied in Jordan & Dimakis, 2020. Our analysis is based on the classical underlying piecewise-affine structure of ReLU networks: on any strict ReLU cell, the logit map $\theta(\cdot)$ is affine with a constant input-logit Jacobian $J_{D}$. Our contribution is to leverage this structure for WDRO, which requires combining the Lipschitz constant of the logit map and the sensitivity of the softmax cross-entropy, or the DLR loss. Our first theoretical result yields an upper bound of WDRO induced by $\bm{L}\triangleq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}$, where $J_{\bm{D}}$ is general Jacobian of the logit map. (See 3.1 for precise definition of $J_{\bm{D}}$.) In addition, we derive a lower bound of WDRO by constructing a concrete and finite worst-case distribution. (See equation 16 for the explicit formulation.) This worst-case distribution is constructed by perturbing the empirical sample along the direction which the logit map is most varied. Moreover, we provide a sufficient condition where our lower and upper bounds match, and simulate an instance to illustrate this tightness, see Figure 2.
@@
    We further analyze the upper and lower bounds of the Wasserstein Distributional Robust Optimization (WDRO) problem for a class of MLP with smooth activation and cross-entropy loss. Unlike ReLU activation or DLR loss, which might create degeneration edges, the chain rule is readily applied in this case, and the Lipschitz constant of the loss is naturally computed by estimating its gradients. Similar to the analysis of the ReLU networks, we obtain the upper bound of the WDRO as $\bm{L}\triangleq 2^{1/s}\max_{x\in\mathcal{X}}\left\|\nabla\theta(x)\right\|_{{r}\to{s}}$ while the worst-case distribution and lower bound are constructed similar to the ReLU networks.
@@
    Finally, we bridge the gap between WDRO theory and adversarial evaluation by introducing the Wasserstein Distributional Attack (WDA), which directly constructs adversarial distributions within the Wasserstein ball rather than restricting to point-wise perturbations. Unlike existing attacks that place all adversarial examples on the $\epsilon$-ball boundary, WDA flexibly interpolates between point-wise ($\kappa=1$) and truly distributional attacks ($\kappa>1$) by supporting adversarial distributions on $2N$ points. This offers a complementary perspective to strong baselines such as AutoAttack and the RobustBench leaderboard (Croce & Hein, 2020; Croce et al., 2021). Empirically, WDA with $\kappa=2$ consistently finds stronger adversarial examples than state-of-the-art methods across diverse settings: achieving lower robust accuracy than APGD-DLR on CIFAR-10/100 with WideResNet backbone (Zagoruyko & Komodakis, 2016) on both $\ell_{\infty}$ and $\ell_{2}$ perturbations. When integrated into the Adaptive Auto Attack framework, our method matches or exceeds the ensemble performance of A3. These results demonstrate that the distributional perspective not only provides tighter theoretical certificates but also yields more effective attacks, validating our claim that existing robustness evaluations underestimate vulnerability by restricting to $\Omega_{\infty}$ rather than the larger $\Omega_{1}$ ambiguity set assumed by certificates.
@@ 2 Preliminaries
Notations
We denote basis vector as $\bm{e}_{k}$; indicator function as $\bm{1}_{\{\cdot\}}$; dirac measure as $\bm{\delta}_{z}$; input dimension $n$, and output dimension as $K$. An empirical dataset is denoted $\{Z^{(1)},\dots,Z^{(N)}\}$ with $Z=(x,y)\in\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{Y}\subset\mathbb{R}^{K}$; empirical distribution $\mathbb{P}_{N}=\sum_{i}\mu_{i}\bm{\delta}_{Z^{(i)}}$ with $Z^{(i)}=(x^{(i)},y^{(i)}=\bm{e}_{k_{i}})$. Norms $\|\cdot\|_{r}$ and $\|\cdot\|_{s}$ are dual with 1/r+1/s=11/{r}+1/{s}=1. For a matrix $A$, $\|A\|_{{r}\to{s}}=\sup_{\|u\|_{r}=1}\|Au\|_{s}$. Rectifier $\cdot]_{+}=\max{\{0,\cdot\}}$. Recession cone $\mathrm{rec}(\cdot)$. Interior set $\mathrm{int}(\cdot)$. Ground cost $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$. Cross-entropy loss $\ell(x,y;\theta)=-\sum_{k=1}^{K}y_{k}\log\operatorname{softmax}(\theta(x))_{k}$. Analogous DLR loss as defined in Croce & Hein (2020). We define the dual-norm maximizer $\mathcal{M}_{r}$ by
@@
and projection $\Pi_{r,{x},\kappa\epsilon}$ by
$\Pi_{r,{x},R}\colon x\mapsto\arg\min_{\xi}\left\{\|\xi-x\|_{2}^{2}\mid\|\xi\|_{r}\leq R\right\}.$ (2)
Wasserstein Distributionally Robust Optimization (WDRO)
Robustness guarantees and certificates aim to make model predictions trustworthy under adversarial manipulation (Wong & Kolter, 2018; Cohen et al., 2019; Salman et al., 2019). Empirical risk minimization model $\inf_{\theta}\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]$ optimizes average performance on the observed data but offers no protection against worst-case shifts nearby. Distributionally robust optimization (DRO) addresses this by choosing parameters that perform well against all distributions within a prescribed neighborhood: $\inf_{\theta}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)].$ Here, the worst-case loss is taken over all admissible distributions $\mathbb{P}\in\mathcal{P}$. The ambiguity (or uncertainty) set $\mathcal{P}$ is often constructed by collecting all distributions $\mathbb{P}$ that are similar to the empirical distribution $\mathbb{P}_{N}$.
In this work, we focus on the Wasserstein ambiguity set, which is a ball centered at $\mathbb{P}_{N}$ under the Wasserstein distance. Given a ground cost $d$ on the space of data $\mathcal{Z}$, the Wasserstein distance (Villani, 2008) between two distributions $\mathbb{P}$ and $\mathbb{Q}$ it is defined as $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\left(\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{Z}\times\mathcal{Z}}d^{p}(z^{\prime},z)\,\mathrm{d}\pi(z^{\prime},z)\right)^{1/p}\text{ for}\,p\in[1,\infty);$, and $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\operatorname{ess.sup}_{{\pi}}(d)\text{ for}\,p=\infty$. Intuitively, the Wasserstein distance between two distributions $\mathbb{P}$ and $\mathbb{Q}$ is defined as the minimum cost to transport the mass of $\mathbb{P}$ to $\mathbb{Q}$. The WDRO problem with a given budget of perturbation $\epsilon>0$ can be written as
@@
It is worth noting that $\mathcal{W}_{d,p}\leq\mathcal{W}_{d,p^{\prime}}$ if $p\geq p^{\prime}$, thus $\Omega_{1}\supseteq\Omega_{p}\supseteq\Omega_{p^{\prime}}\supseteq\Omega_{\infty}$ (see Figure 1).
Lipschitz Certificate
@@
where $L$ is any Lipschitz constant of $z\mapsto\ell(z;\theta)$ with respect to the ground cost. This inequality follows from weak duality and is widely used to make the WDRO objective tractable: one replaces the inner maximization by the surrogate $L\epsilon$ and then controls $L$ (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024). In practice, estimating $L$ reduces to bounding the network’s (global or local) Lipschitz modulus, e.g., fast global products of per-layer operator norms (Virmaux & Scaman, 2018) or tighter activation-aware/local certificates (Jordan & Dimakis, 2020; Shi et al., 2022).
@@
Adversarial attack methods often construct a perturbed distribution by shifting each sample $X^{(i)}$ along a specific adversarial direction $u^{(i)}$ to get $X_{\mathrm{adv}}^{(i)}$ (Goodfellow et al., 2014; Moosavi-Dezfooli et al., 2016; Carlini & Wagner, 2017). These methods are essentially point-wise attacks, which draws a distribution $\mathbb{P}_{\mathrm{adv}}=\sum_{i=1}^{N}\frac{1}{N}\bm{\delta}_{X^{(i)}_{\mathrm{adv}}}$ in the Wasserstein ambiguity set $\Omega_{p}=\{\mathbb{P}\colon\mathcal{W}_{d,p}\leq\epsilon\}$ when $p=\infty$ (see Figure 1). Whereas, in the $p=1$ case, the ambiguity set only constrains the average transportation cost under an optimal coupling. Hence, the adversary may move some points farther and others less as long as the mean cost stays within budget. This creates a significant gap between the robustness measured against $\Omega_{p=\infty}$ attacks and the theoretical robustness or Lipschitz certificates 4 which are developed for $\Omega_{p=1}$ (Mohajerin Esfahani & Kuhn, 2018; Carlini et al., 2019; Rice et al., 2021).
@@ 3 Tractable Interpretation of WDRO for Neural Networks
For certain shallow and convex models (e.g., linear regression, support vector machines, etc.), the tractable representation of the WDRO problem 3 is well-established in the literature (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024). This tractable form enables a computational advantage and provides a clear interpretation of the robustness of regularization mechanisms. In that line of work, the Lipschitz constant often provides a practical and tight upper bound of the corresponding upper bounds. However, when the loss is non-convex, the Lipschitz certificate is not always tight, as outlined in the following remark.
@@
Consider a single-point empirical $\mathbb{P}_{N}=\bm{\delta}_{\{X^{(1)}=2\}}$ and a loss given by
@@
Then $\ell$ is Lipschitz with modulus 1, however $\sup_{\Omega_{1}}\mathbb{E}_{\mathbb{P}}[\ell(X)]=\mathbb{E}_{\mathbb{P}_{N}}[\ell(X)]+\frac{\epsilon}{2}\quad\text{for any}\,\epsilon>0.$
As presented in the following sections, our main theoretical results (Theorem 3.1 and 3.3) show that Lipschitz modulus provides a tight upper bound for the WDRO problem 3 for a class of ReLU neural networks and smooth activated neural networks.

### v2

@@ Abstract
Wasserstein distributionally robust optimization (WDRO) provides a framework for adversarial robustness, yet existing methods based on global Lipschitz continuity or strong duality often yield loose upper bounds or require prohibitive computation. We address these limitations with a primal approach and adopt a notion of exact Lipschitz certificates to tighten this upper bound of WDRO. For ReLU networks, we leverage the piecewise-affine structure on activation cells to obtain an exact tractable characterization of the corresponding WDRO problem. We further extend our analysis to modern architectures with smooth activations (e.g., GELU, SiLU), such as Transformers. Additionally, we propose novel Wasserstein Distributional Attacks (WDA, WDA++) that construct candidates for the worst-case distribution. Compared to existing attacks that are restricted to point-wise perturbations, our methods offer greater flexibility in the number and location of attack points. Extensive evaluations demonstrate that our proposed framework achieves competitive robust accuracy against state-of-the-art baselines while offering tighter certificates than existing methods. Our code is available at https://github.com/OLab-Repo/WDA
@@ 1 Introduction
Modern deep networks achieve remarkable accuracy yet remain fragile to distribution shift and adversarial perturbations Szegedy et al. (2014); Goodfellow et al. (2014); Kurakin et al. (2018); Hendrycks and Dietterich (2019); Ovadia et al. (2019); Taori et al. (2020); Koh et al. (2021), raising concerns about their reliability in deployment. A principled avenue for robustness is Wasserstein distributionally robust optimization (WDRO, Mohajerin Esfahani and Kuhn 2018; Gao and Kleywegt 2023), which controls worst-case test risk over an ambiguity set within a Wasserstein ball around the empirical distribution and admits tight dual characterizations from optimal transport (Villani and others, 2008; Santambrogio, 2015). While numerous defences have been proposed, a fundamental gap persists between theoretical robustness certificates and practical adversarial evaluation: existing Lipschitz-based certificates often provide loose upper bounds that vastly overestimate the true worst-case loss (Virmaux and Scaman, 2018), while standard attacks restrict perturbations to fixed-radius balls around individual points (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018). This mismatch stems from two limitations: certificates typically rely on global worst-case analysis that ignores the actual network geometry traversed by data, and attacks consider only point-wise perturbations rather than distributional shifts permitted by Wasserstein threat models (Singh et al., 2018; Gao and Kleywegt, 2023). The discrepancy is particularly pronounced for modern architectures with ReLU activations, where the network behaves as a piecewise-affine function whose local properties vary dramatically across regions (Jordan and Dimakis, 2020), and those with smooth activations (GELU, SiLU/Swish) exhibit complex nonlinear geometry (Hendrycks and Gimpel, 2016; Ramachandran et al., 2017; Elfwing et al., 2018). In this work, we aim to address both sides of this gap: our contributions can be summarized as follows.
@@
    For a class of networks with Rectified Linear Unit (ReLU) activations (Nair and Hinton, 2010), we analyse the upper and lower bounds of the Wasserstein Distributional Robust Optimization (WDRO) problem by connecting with the tight Lipschitz constant studied in Jordan and Dimakis, 2020. Our first theoretical result yields an upper bound of WDRO induced by $\bm{L}\triangleq 2^{1/r}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}$, where $J_{\bm{D}}$ is general Jacobian of the logit map, see 3.1. In addition, we derive a lower bound of WDRO by constructing a concrete and finite worst-case distribution, see equation 17. This worst-case distribution is constructed by perturbing the empirical sample along the direction in which the logit map is most varied. Moreover, we provide a sufficient condition where our lower and upper bounds match, and simulate an instance to illustrate this tightness, see Figure 1. It is worth noting that not every Lipschitz constant yields a tight bound of WDRO, see Remark Remark.
@@
    We further analyse the upper and lower bounds of the WDRO problem for a class of architectures with smooth activation and cross-entropy loss. Unlike ReLU activation or DLR loss, which might create degeneration edges, the global differentiability of smooth networks enables an exact characterization of the Lipschitz constant via the gradient norm. Similar to the analysis of the ReLU networks, we obtain the upper bound of the WDRO as $\bm{L}\triangleq 2^{1/r}\max_{x\in\mathcal{X}}\left\|\nabla\theta(x)\right\|_{{r}\to{s}}$ while the worst-case distribution and lower bound are constructed similar to the ReLU networks.
@@
    Finally, we bridge the gap between WDRO theory and adversarial evaluation by introducing the Wasserstein Distributional Attack (WDA) and its adaptive variant (WDA++), which directly constructs adversarial distributions within the Wasserstein ball rather than restricting to point-wise perturbations. Unlike existing attacks that place all adversarial examples on the $\epsilon$-ball boundary, WDA flexibly interpolates between point-wise ($\kappa=1$) and truly distributional attacks ($\kappa>1$) by supporting adversarial distributions on $2N$ points, see Figure 2. WDA++ further optimizes the $2N$-support distribution via adaptive transport budget allocation. Empirically, WDA and WDA++ consistently find stronger adversarial examples than state-of-the-art methods across diverse settings (Table 1). When integrated into the Adaptive Auto Attack framework (Liu et al., 2022), our method matches or exceeds the ensemble performance of A3. Our results demonstrate that the distributional perspective not only provides tighter theoretical certificates but also yields more effective attacks. Crucially, this exposes the limitations of current evaluation standards (e.g., AutoAttack (Croce and Hein, 2020), RobustBench (Croce et al., 2021)) which significantly overestimate robustness by restricting to $\Omega_{\infty}$ rather than the larger $\Omega_{1}$ ambiguity set assumed by certificates.
Figure 1: WDRO bounds and PGD attack loss for a ReLU classifier with $n=K=2$ and one hidden layer of dimension 8. Lower bound curves are the cumulative $\bm{l}$ as more reachable activation masks are considered.
@@ 2 Preliminaries
Notations.
We denote basis vector as $\bm{e}_{k}$; indicator function as $\bm{1}_{\{\cdot\}}$; Dirac measure as $\bm{\delta}_{z}$; input dimension $n$, and output dimension as $K$. An empirical dataset is denoted $\{Z^{(1)},\dots,Z^{(N)}\}$ with $Z=(x,y)\in\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{Y}\subset\mathbb{R}^{K}$; empirical distribution $\mathbb{P}_{N}=\sum_{i}\mu_{i}\bm{\delta}_{Z^{(i)}}$ with $Z^{(i)}=(x^{(i)},y^{(i)}=\bm{e}_{k_{i}})$. Norms $\|\cdot\|_{r}$ and $\|\cdot\|_{s}$ are dual with $1/r+1/s=1$. For a matrix $A$, $\|A\|_{{r}\to{s}}=\sup_{\|u\|_{r}=1}\|Au\|_{s}$. Rectifier $(\cdot)_{+}=\max{\{0,\cdot\}}$. Recession cone $\mathrm{rec}(\cdot)$. Interior set $\mathrm{int}(\cdot)$. Ground cost $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$. Cross-entropy loss $\ell(x,y;\theta)=-\sum_{k=1}^{K}y_{k}\log\operatorname{softmax}(\theta(x))_{k}$. Analogous DLR loss as defined in Croce and Hein (2020). Dual-norm maximizer $\mathcal{M}_{r}$
@@
and projector $\Pi_{r,{x},R}$
[…]
WDRO.
Robustness certificates aim to make model predictions trustworthy under adversarial manipulation (Wong and Kolter, 2018; Cohen et al., 2019; Salman et al., 2019). The empirical risk minimization model $\inf_{\theta}\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]$ optimizes average performance on the observed data but offers no protection against worst-case shifts nearby. Distributionally robust optimization (DRO) addresses this by choosing parameters that perform well against all distributions within a prescribed neighbourhood: $\inf_{\theta}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)].$ Here, the worst-case loss is taken over all admissible distributions $\mathbb{P}\in\mathcal{P}$. The ambiguity (or uncertainty) set $\mathcal{P}$ is often constructed by collecting all distributions $\mathbb{P}$ that are similar to the empirical distribution $\mathbb{P}_{N}$.
In this work, we focus on the Wasserstein ambiguity set, which is a ball centred at $\mathbb{P}_{N}$ under the Wasserstein distance. Given a ground cost $d$ on the space of data $\mathcal{Z}$, the Wasserstein distance $\mathcal{W}_{d,p}$ (Villani and others, 2008) between two distributions $\mathbb{P}$ and $\mathbb{Q}$ is defined for $p\in[1,\infty)$ as
$\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\left(\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{Z}\times\mathcal{Z}}d^{p}(z^{\prime},z)\,\mathrm{d}\pi(z^{\prime},z)\right)^{1/p}$
and $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\operatorname{ess.sup}_{{\pi}}(d)$ for $p=\infty$. Intuitively, the Wasserstein distance between two distributions $\mathbb{P}$ and $\mathbb{Q}$ is defined as the minimum cost to transport the mass of $\mathbb{P}$ to $\mathbb{Q}$. The WDRO problem with a given budget of perturbation $\epsilon>0$ can be written as
@@
It is worth noting that $\mathcal{W}_{d,p}\leq\mathcal{W}_{d,p^{\prime}}$ if $p\leq p^{\prime}$, thus $\Omega_{1}\supseteq\Omega_{p}\supseteq\Omega_{p^{\prime}}\supseteq\Omega_{\infty}$ (see Figure 2). For more details on Wasserstein distributionally robust optimization, we refer reader to Kuhn et al. (2019) and our Appendix A.
Lipschitz Certificate.
@@
where $L$ is any Lipschitz constant of $z\mapsto\ell(z;\theta)$ with respect to the ground cost. This inequality follows from weak duality and is widely used to make the WDRO objective tractable: one replaces the inner maximization by the surrogate $L\epsilon$ and then controls $L$ (Mohajerin Esfahani and Kuhn, 2018; Blanchet et al., 2019; Gao and Kleywegt, 2023; Gao et al., 2024). In practice, estimating $L$ reduces to bounding the network’s (global or local) Lipschitz modulus, e.g., fast global products of per-layer operator norms (Virmaux and Scaman, 2018) or tighter activation-aware/local certificates (Jordan and Dimakis, 2020; Shi et al., 2022).
@@
Adversarial attack methods often construct a perturbed distribution by shifting each sample $X^{(i)}$ along a specific adversarial direction $u^{(i)}$ to get $X_{\mathrm{adv}}^{(i)}$ (Goodfellow et al., 2014; Moosavi-Dezfooli et al., 2016; Carlini and Wagner, 2017). These methods are essentially point-wise attacks, which draws a distribution $\mathbb{P}_{\mathrm{adv}}={\scriptstyle\sum_{i=1}^{N}\frac{1}{N}}\bm{\delta}_{(X^{(i)}_{\mathrm{adv}},Y^{(i)})}$ in the Wasserstein ambiguity set $\Omega_{p}=\{\mathbb{P}\colon\mathcal{W}_{d,p}\leq\epsilon\}$ when $p=\infty$ (see Figure 2). Whereas, in the $p=1$ case, the ambiguity set only constrains the average transportation cost under an optimal coupling. Hence, the adversary may move some points farther and others less as long as the mean cost stays within budget. This creates a significant gap between the robustness measured against $\Omega_{p=\infty}$ attacks and the theoretical robustness or Lipschitz certificates equation 4 which are developed for $\Omega_{p=1}$ (Mohajerin Esfahani and Kuhn, 2018; Carlini et al., 2019; Rice et al., 2021).
@@ 3 Tractable Interpretation of WDRO for Neural Networks
For certain shallow and convex models (e.g., linear regression, support vector machines, etc.), the tractable representation of the WDRO problem equation 3 is well-established in the literature (Mohajerin Esfahani and Kuhn, 2018; Blanchet et al., 2019; Gao and Kleywegt, 2023; Gao et al., 2024). This tractable form enables a computational advantage and provides a clear interpretation of the robustness of regularization mechanisms. In that line of work, the Lipschitz constant often provides a practical and tight upper bound of the corresponding upper bounds. However, when the loss is non-convex, the Lipschitz certificate is not always tight, as outlined in the following remark.
@@
Consider a single-point empirical distribution $\mathbb{P}_{N}=\bm{\delta}_{\{X^{(1)}=2\}}$ and a loss function given by
@@
Then $\ell$ is Lipschitz with Lipschitz constant 1, however for any $\epsilon>0$,
$\sup_{\Omega_{1}}\mathbb{E}_{\mathbb{P}}[\ell(X)]=\mathbb{E}_{\mathbb{P}_{N}}[\ell(X)]+\frac{1}{2}\epsilon.$
In the following, our main theoretical results (Theorems 3.1 and 3.3) show that Lipschitz modulus provides a tight upper bound for the WDRO problem equation 3 for a class of ReLU neural networks and smooth activated neural networks.